Consider the following two mutually exclusive projects:

  

Year Cash Flow (A) Cash Flow (B)
0 –$230,324        –$16,246         
1 27,000        5,466         
2 59,000        8,622         
3 56,000        13,991         
4 426,000        9,861         

  

Whichever project you choose, if any, you require a 6 percent return on your investment.
d. What is the discounted payback period for Project B?
e. What is the NPV for Project A?

  

g. What is the IRR for Project A?
h. What is the IRR for Project B?
i. What is the profitability index for Project A?
j. What is the profitability index for Project B?
Answer #1

к LM arc D8 fx =SUM(D3:07) - A 1 Project A 2 year Cash flows pv@6% Present value Cumulative cash flows Cumulative Discounted

f =SUM(D3:07) B C - A 1 Project A 2 year 30 4 1 pv@6% Cash flows -230324 27000 59000 56000 426000 =IRR(B3:B7) DWN =C3/1.06 =C